Write an equation with slope 3 through (6,-5)
Through (-4,-6) and (-5,8)
Through (6,-1) and parallel to y=3x+4

Answer by lwsshak3(11628)

Equation of a line: y=mx+b, m=slope, b=y-intercept
Write an equation with slope 3 through (6,-5)
y=3x+b
solve for b using coordinates of given point (6,-5)
-5=3*6+b
b=-5-18=-23
Equation:y=3x-23
..
Through (-4,-6) and (-5,8)
slope=∆y/∆x=(8-(-6))/(-5-(-4))=14/-1=-14
Equation: y=-14x+b
-6=-14*-4+b
b=-6-56=-62
Equation: y=-14x-62
..
Through (6,-1) and parallel to y=3x+4
slope=3
y=3x+b
-1=3*6+b
b=-19
equation: y=3x-19